# Link Representation Learning for Probabilistic Travel Time Estimation

### v1

@@ V-C Interpretability Analysis of Link Embedding Vectors
We analyze the interpretability of link embedding vectors and correlations on the Harbin dataset through data visualization. We visualized the learned link correlations $\Sigma_{d}$ and $\Sigma_{p}$ in two time periods (9:00-10:00 AM and 9:00-10:00 PM), then compared them with the corresponding 2-hop real adjacency matrix, as shown in Figure 4(a,b,c,d,e). Despite not incorporating any network geometry information in the generation of link representations $L$ and $H$, the derived link correlations still effectively capture the essential geometry information of the road network. Compared to the real adjacency matrix, the learned link correlations are more flexible and can adaptively construct correlations for different links. For example, some link correlations may have weak extensions, with high correlations lasting only 1 or 2 hops, while other links may exhibit stronger extensions, with high correlations persisting over multiple hops. In $\Sigma_{d}$, the learned correlations are smoother, reflecting global low-frequency characteristics, while in $\Sigma_{p}$, the learned correlations are sharper, representing high-frequency characteristics. Additionally, the correlation of links shows noticeable variations across different periods: it is stronger during peak hours and relatively weaker at night. In Figure 4(f,g), we randomly selected a link and visualized its correlations with other links in 9:00-10:00 AM on the real map and we can see that $\Sigma_{d}$ indeed demonstrate long-range correlations, while the correlation structure in $\Sigma_{p}$ is more local. Overall, the obtained $\Sigma_{d}$ and $\Sigma_{p}$ are consistent with our prior specification in Eq. (2), where $\Sigma_{d}$ and $\Sigma_{p}$ are used for modeling inter-trip correlation and intra-trip correlation, respectively. Next, we utilized Principal Component Analysis (PCA) to project the comprehensive representations of the link vectors $\{L,H\}$ into a two-dimensional space, and compared them with their actual map locations. As illustrated in Figure 4(h,i), vectors that are nearer in the representation space generally correspond to links that are closer on the map. Moreover, link representation vectors that run in the same direction on the same road tend to have higher similarity, whereas vectors from different roads show lower similarity. These results further confirm the effectiveness of $\{L,H\}$ as link representations.
@@
Figure 5: Cumulative travel time (min) estimation: shaded area shows $\mu\pm\sigma$.
@@
We next visualize the mean and variance of travel times at the link level from one selected trip, which is divided into 8 segments with equal intervals of GPS points. Two scenarios are considered: with condition and without condition. In the conditional case, the travel times of 32 completed trips are used as conditional information to adjust the distribution of the travel time for each link in the query trip. In the without-conditional case, only the learned link representation vectors are used to compute the mean and variance of the link travel times. The results, as shown in Figure 5, demonstrate that our model maintains good estimation performance and stability at the link level. The introduction of conditional information significantly reduces the prediction error and variance, and this positive adjustment becomes more apparent as the trip progresses.
(a) The batch size of joint distribution.
(b) The granularity of subsampling.
(c) The granularity of time discretization.
Figure 6: Performance metrics associated with different configurations.
@@ V-D1 The batch size of the joint distribution
We conducted a performance comparison of our model across various batch sizes of the joint distribution. To isolate the effects of other factors, we avoided introducing conditional trips, data augmentation, and time discretization. Therefore, our comparison solely reflects the performance achieved through multi-trip joint modeling using the original data. To enhance experimental efficiency, we adopted an exponential increase in the batch size, evaluating performance using {1, 2, 4, 8, 16, 32, 64, 128, 256}. As shown in Figure 6(a), the multi-trip joint modeling significantly improved prediction performance. When the batch size in the joint modeling increased, the MAPE saw a significant decline. However, upon reaching a specific threshold (approximately 32 trips), this decreasing trend slowed, and with additional increases in the number of trips, the MAPE started to increase marginally. We consider that multi-trip joint modeling offers additional information, allowing the model to identify relationships between trips. However, an excessively large batch size might strain the model’s optimization, making it more difficult to further improve performance and potentially leading to a performance plateau.
@@ V-D3 The granularity of time discretization
We conducted a performance evaluation on temporal discretization based on the parameters determined in the preliminary experiments. We divided one day into $p=\{1,2,3,4,6,8,12,24,72\}$ periods and independently learned link embedding vectors for each period using separate parameters. As shown in Figure 6(c), the results indicate that the introduction of new trainable parameters improved the model performance. Fine-grained time discretization enables more accurate learning of link correlations, thereby enhancing model performance.
@@ V-E Analysis of Time Complexity
The time complexity of ProbTTE mainly comes from the mean and diagonal mapping, correlation computation, maximum likelihood estimation, and conditional distribution estimation. The time complexity of mean and diagonal mapping is $O(Nr)$. In correlation computation, the computational complexity is influenced by the number of trips in one joint probability distribution (batch size) $b$. $N$ samples are divided into $N/b$ batches, and the correlation of trips is calculated within each batch. The time complexity for calculating the inner product pairwise within each batch is $O(b^{2}r)$. So the total time complexity for correlation computation is $N/b*O(b^{2}r)=O(Nbr)$, where $r\ll N$. The final part of the computational complexity arises from maximum likelihood estimation and conditional distribution estimation. In these processes, with the help of the Woodbury Matrix Identity and the Matrix Determinant Lemma, the time complexity for inverting the covariance matrix and calculating the determinant is $O(rb^{2})$, for $N/b$ batches, the total time complexity is $O(Nbr)$. Therefore, the overall time complexity of ProbTTE is $O(Nr)+O(Nbr)+O(Nbr)=O(Nbr)$, where $b\ll N$, and $r\ll b$ in general.
We also conducted a comparison of the actual runtime of our model and the baselines. We recorded the time it took to train and infer one epoch for each model on the Chengdu and Harbin datasets. The device used for execution was an NVIDIA Tesla V100 GPU, and the runtime is shown in Table IV. The results indicate that our model achieves excellent performance without introducing excessive computational overhead.
TABLE IV: Runtime (s) of models.
Model DeepTTE HierETA MulT-TTE DeepGTT GMDNet ProbTTE
@@ VI Conclusion
In this paper, we propose ProbTTE, a probabilistic TTE model designed to capture the joint probability distribution of multiple trips. By introducing a covariance representation paradigm based on low-rank link representation, we effectively modeled correlations between trips with low time complexity and the learned representation gives favorable contraction properties. The use of subsampling data augmentation facilitated gradient propagation at the link level, resulting in the interpretable optimization of link representation vectors. ProbTTE demonstrated exceptional performance on real-world datasets, outperforming state-of-the-art deterministic and probabilistic baselines by 12.11% and 13.34% on average, respectively. Ablation experiments and visual analyses confirmed the effectiveness of each model component and the interpretability of the learned link representation vectors. Additionally, time complexity analysis and runtime comparisons illustrated that ProbTTE maintains a low computational burden while achieving superior performance in jointly modeling multiple trips.
In terms of future work, we plan to further explore the continuous modeling of link embedding vectors. Specifically, dynamic graph neural networks can be designed to learn low-rank temporal feature representations on a graph Laplacian basis from observed trips. Recurrent neural networks can subsequently be leveraged to capture the evolution of link embedding vectors based on the contextual information of observed trips. We believe that these enhancements will enable the model to better perceive temporal trends and facilitate the modeling of travel time distributions across multiple time horizons.

### v2

@@ V-C Interpretability Analysis of Link Embedding Vectors
Figure 4: Link correlation visualization. (a). Learned inter-trip link correlation during 9:00-10:$0$ AM. (b). Learned intra-trip link correlation during 9:00-10:$0$ AM. (c). Learned inter-trip link correlation during 9:00-10:$0$ PM. (d). Learned intra-trip link correlation during 9:00-10:$0$ PM. (e). Real adjacency matrix. (f). Link correlation heatmap from inter-trip. (g). Link correlation heatmap from intra-trip. (h). Visualization of embedding vectors by Principal Component Analysis (PCA). (i). Visualization of real link locations
@@
We analyze the interpretability of link embedding vectors and correlations on the Harbin dataset through data visualization. We visualized the learned inter-trip (from covariance matrix $\Sigma_{d}$) and intra-trip (from covariance matrix $\Sigma_{p}$) link correlations in two time periods (9:00-10:00 AM and 9:00-10:00 PM), then compared them with the corresponding 2-hop real adjacency matrix, as shown in Figure 4(a,b,c,d,e). Despite not incorporating any network geometry information in the generation of link representations $L$ and $H$, the derived link correlations still effectively capture the essential geometry information of the road network. Compared to the real adjacency matrix, the learned link correlations are more flexible and can adaptively construct correlations for different links. For example, some link correlations may have weak extensions, with high correlations lasting only 1 or 2 hops, while other links may exhibit stronger extensions, with high correlations persisting over multiple hops. The learned inter-trip correlations are smoother, reflecting global low-frequency characteristics, while the learned intra-trip correlations are sharper, representing high-frequency characteristics. Additionally, the correlation of links shows noticeable variations across different periods: it is stronger during peak hours and relatively weaker at night. In Figure 4(f,g), we randomly selected a link and visualized its correlations with other links in 9:00-10:00 AM on the real map and we can see that inter-trip correlation indeed demonstrates broad-scale correlations, while the correlation structure from intra-trip is more local. Additionally, intra-trip correlations tend to be stronger between links on the same road, such as adjacent segments along a main road, compared to links on different roads, such as those connecting a main road to its adjacent side streets. This also indicates that features associated with inter-trip correlations tend to be smoother, whereas those related to intra-trip correlations are sharper. Overall, the obtained $\Sigma_{d}$ and $\Sigma_{p}$ are consistent with our prior specification in Eq. (2).
Next, we utilized Principal Component Analysis (PCA) to project the comprehensive representations of the link vectors $\{L,H\}$ into a two-dimensional space, and compared them with their actual map locations. As illustrated in Figure 4(h,i), vectors that are nearer in the representation space generally correspond to links that are closer on the map. Moreover, link representation vectors that run in the same direction on the same road tend to have higher similarity, whereas vectors from different roads show lower similarity. These results further confirm the effectiveness of $\{L,H\}$ as link representations.
@@
Figure 5: Cumulative travel time (min) estimation: shaded area shows $\mu\pm\sigma$. (a). Trip sample from Chengdu. (b). Trip sample from Harbin.
@@
We also visualize the mean and variance of travel times at the link level from selected trips from Chengdu and Harbin, which are divided into several segments with equal intervals of GPS points. Two scenarios are considered: with condition and without condition. In the conditional case, the travel times of 32 completed trips are used as conditional information to adjust the distribution of the travel time for each link in the query trip. In the without-conditional case, only the learned link representation vectors are used to compute the mean and variance of the link travel times. The results, as shown in Figure 5(a,b), demonstrate that our model maintains good estimation performance and stability at the link level. The introduction of conditional information significantly reduces the prediction error and variance, and this positive adjustment becomes more apparent as the trip progresses.
Figure 6: Performance metrics associated with different configurations. (a). The batch size of joint distribution. (b). The granularity of subsampling. (c). The granularity of time discretization. (d). The rank of representation matrix.
@@ V-D Analysis of Model Parameters
Regarding the important parameters in the model, including the number of trips in the joint distribution, the amount of data augmentation, the level of temporal discretization, the value of rank, and the hyperparameter $\alpha$, we tested the model’s performance under different values to determine the optimal settings for these parameters.
@@ V-D1 The batch size of the joint distribution
We conducted a performance comparison of our model across various batch sizes of the joint distribution. To isolate the effects of other factors, we avoided introducing conditional trips, data augmentation, and time discretization. Therefore, our comparison solely reflects the performance achieved through multi-trip joint modeling using the original data. To enhance experimental efficiency, we adopted an exponential increase in the batch size, evaluating performance using {1, 2, 4, 8, 16, 32, 64, 128, 256}. As shown in Figure 6(a), the multi-trip joint modeling significantly improved prediction performance. When the batch size in the joint modeling increased, the MAPE and CRPS saw a significant decline. However, upon reaching a specific threshold (approximately 32 trips), this decreasing trend slowed, and with additional increases in the number of trips, the MAPE and CRPS started to increase marginally. We consider that multi-trip joint modeling offers additional information, allowing the model to identify relationships between trips. However, an excessively large batch size might strain the model’s optimization, making it more difficult to further improve performance and potentially leading to a performance plateau.
@@ V-D3 The granularity of time discretization
We conducted a performance evaluation on temporal discretization based on the parameters determined in the preliminary experiments. We divided one day into $p=\{1,2,3,4,6,8,12,24,72\}$ periods and independently learned link embedding vectors for each period using separate parameters. As shown in Figure 6(c), the results indicate that the introduction of new trainable parameters improved the model performance. Fine-grained time discretization enables more accurate learning of link correlations, thereby enhancing model performance. However, overly fine temporal discretization can result in an insufficient number of samples for each time period, leading to excessively sparse link coverage, which may cause the model to overfit and ultimately degrade its performance.
V-D4 The rank of representation matrix
Based on the parameters determined from the above experiments, we further evaluated the model performance under different ranks {4, 8, 16, 32, 64, 128}, as shown in Figure 6(d). When the rank is small, both MAPE and CRPS metrics are significantly impacted. As the rank increases, MAPE reaches optimal performance first, while CRPS requires a higher rank to achieve its optimal performance. This indicates that the correlation between links (the covariance of the joint distribution) may be more challenging to learn than the mean. Considering that both MAPE and CRPS achieve relatively optimal performance without introducing an excessive number of parameters, we ultimately selected a rank of 32.
V-D5 The weight of orthogonal regularization
We also tested the effect of orthogonal regularization on model training. We compared the training convergence process under three conditions: without orthogonal regularization ($\alpha=0$), with an appropriate regularization weight ($\alpha=0.2$), and with a larger regularization weight ($\alpha=1$). The results are shown in Figure 7. When no orthogonal regularization is applied ($\alpha=0$), the model converges at around the 74th epoch. With the introduction of orthogonal regularization, the convergence speed improves. When $\alpha=0.2$, the model converges at the 65th epoch, representing a 12.16% increase in convergence speed. When $\alpha=1$, the model converges at the 53rd epoch, achieving a 28.38% increase in convergence speed. However, an excessively high regularization weight can limit the model’s expressive capacity, resulting in slight increases in both MAPE and CRPS. Therefore, we ultimately selected $\alpha=0.2$ as the orthogonal regularization weight for the model.
Figure 7: Training process under different weighted orthogonal regularization.
@@ V-E Analysis of Time Complexity
TABLE V: Runtime (s) of models.
Model DeepTTE HierETA MulT-TTE DeepGTT GMDNet ProbETA
@@
The time complexity of ProbETA mainly comes from the mean and diagonal mapping, correlation computation, maximum likelihood estimation, and conditional distribution estimation. The time complexity of mean and diagonal mapping is $O(Nr)$. In correlation computation, the computational complexity is influenced by the number of trips in one joint probability distribution (batch size) $b$. NN samples are divided into $N/b$ batches, and the correlation of trips is calculated within each batch. The time complexity for calculating the inner product pairwise within each batch is $O(b^{2}r)$. So the total time complexity for correlation computation is $N/b*O(b^{2}r)=O(Nbr)$, where $r\ll N$. The final part of the computational complexity arises from maximum likelihood estimation and conditional distribution estimation. In these processes, with the help of the Woodbury Matrix Identity and the Matrix Determinant Lemma, the time complexity for inverting the covariance matrix and calculating the determinant is $O(rb^{2})$, for $N/b$ batches, the total time complexity is $O(Nbr)$. Therefore, the overall time complexity of ProbETA is $O(Nr)+O(Nbr)+O(Nbr)=O(Nbr)$, where $b\ll N$, and $r\ll b$ in general.
We also conducted a comparison of the actual runtime of our model against the baseline models. We recorded the time required to train and infer one epoch for each model on the Chengdu and Harbin datasets. The execution was performed on an NVIDIA Tesla V100 GPU, and the runtime is presented in Table V. Compared to some baseline models, our model exhibits a slight increase in training time; however, we consider this to be acceptable, as we have modeled the intricate joint distribution of trips/links, including day-specific and trip-specific random effects. This enhancement effectively improves the model’s performance, while the training time remains within the same order of magnitude as the baselines.
@@ VI Conclusion
In this paper, we propose ProbETA, a probabilistic TTE model designed to capture the joint probability distribution of multiple trips. By introducing a covariance representation paradigm based on low-rank link representation, we effectively modeled correlations between trips with low time complexity and the learned representation gives favorable contraction properties. The use of subsampling data augmentation facilitated gradient propagation at the link level, resulting in the interpretable optimization of link representation vectors. ProbETA demonstrated exceptional performance on real-world datasets, outperforming state-of-the-art deterministic and probabilistic baselines by 12.11% and 13.34% on average, respectively. Ablation experiments and visual analyses confirmed the effectiveness of each model component and the interpretability of the learned link representation vectors. Additionally, time complexity analysis and runtime comparisons illustrated that ProbETA maintains a low computational burden while achieving superior performance in jointly modeling multiple trips.
In terms of future work, we plan to further explore the continuous modeling of link embedding vectors. Specifically, dynamic graph neural networks can be designed to learn low-rank temporal feature representations on a graph Laplacian basis from observed trips. Subsequently, recurrent neural networks can be leveraged to capture the evolution of link embedding vectors based on the contextual information of observed trips. We believe that these enhancements will enable the model to better perceive temporal trends and address the issue of spatiotemporal sparsity in trip data.